JET-COOLED ELECTRONIC SPECTROSCOPY OF ZrC

ANTHONY J. MERER and JAMES R.D. PEERS, Department of Chemistry, University of British Columbia, Vancouver, BC, Canada V6T 1Z1; \underline{SCOTT J. RIXON}, Department of Physics and Astronomy, University of British Columbia, 6224 Agricultural Road, Vancouver, BC, Canada V6T 1Z1.

Zirconium monocarbide is produced in high yield from the reaction of laser- ablated Zr metal with a helium/methane (1%) gas mixture under supersonic free jet conditions. The ground state is X^{1}Σ^{+}, but the low-lying a^{3}Σ^{+} state has also been identified by high-resolution laser-induced fluorescence; this has T_{0} = 690 cm^{−1} and ω_{e} = 885 cm^{−1}. Molecular constants have been derived for ^{90}ZrC (51.45% natural abundance) for both states; the corresponding constants of the minor isotopomers ^{92}ZrC (17.15%) and ^{94}ZrC (17.38%) have also been determined for a^{3}Σ^{+}, and scale isotopically within experimental error. The principal electronic bands in the visible region form a strongly perturbed ^{3}Π_{r} − a^{3}Σ^{+} system, of which the origin band has been analysed in detail. The perturbations in the ^{3}Π_{r} state appear to be caused by singlet states.